Fermi National Laboratory
 

Operation Summary for February 28, 2011 - March 2, 2011

 

Monday Feb. 28

The day shift began with the Tevatron (TEV) recovering from a quench, with Operations monitoring the Recycler stash (194.9E10), with the Antiproton Source (Pbar) stacking (15.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target LCW problems, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations had the TEV recovered by 9:35 AM. Operations put the TEV into HEP shot setup and established store 8540 at 10:33 AM with an average initial luminosity (AIL) of 234.76E30.

During the evening shift, Operations reported a MI electrical safety system (ESS) ground fault at 10:26 PM that tripped off MI, causing the loss of the stash, and halted stacking.

Tuesday March 1

The midnight shift began with Operations monitoring store 8540 (56.77E30) and the Pbar stack (27.1E10), with no stash due to a MI ESS ground fault.

During the midnight shift, Interlock personnel investigated and resolved the ESS problem. Operations reestablished MI and MiniBooNE beam at 12:54 AM. Operators reported that they halted stacking at 6:48 AM in preparation for a Pbar water cage access to switch from one dump LCW system to a parallel system. Operations terminated store 8540 at 7:11 AM and put the TEV into shot setup.

During the day shift, Operations established store 8541 at 8:43 AM with an AIL of 180.27E30. At 10:55 AM, the MI ESS registered another ground fault. The stash was lost. Interlock technicians traced the problem to a door interlock, which they repaired. Operations resumed sending beam to MiniBooNE at 12:52 PM and to Pbar at 1:27 PM.

During the evening shift, Operations monitored the accelerator complex.

Wednesday March 2

The midnight shift began with Operations monitoring store 8541 (39.29E30) and the stash (271.4E10), with Pbar stacking (27.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8541 at 1:03 AM, put the TEV into shot setup, and established store 8542 at 2:23 AM with an AIL of 281.69E30.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TEV personnel to conduct a destructive end-of-store study to move the interaction points for the experiments. CDF and D0 will have to turn off for this study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/02/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory